40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 21 19:28:03 2021, Anchal, Update, PSL, PSL/IFO recovery 
    Reply  Wed Apr 21 19:43:20 2021, Koji, Update, PSL, New HEPA speed control P_20210421_193637.jpgP_20210421_193627.jpg
       Reply  Thu Apr 22 14:41:55 2021, Chub, Update, PSL, New HEPA speed control 40M_PSL_HEPA.jpg
          Reply  Fri Apr 23 18:00:02 2021, gautam, Update, PSL, HEPA speed lowered HEPAdiag.pdf
             Reply  Fri Apr 23 19:26:58 2021, Koji, Update, PSL, HEPA speed lowered 
             Reply  Fri May 14 17:45:05 2021, rana, Update, PSL, HEPA speed raised 
                Reply  Tue May 18 00:52:38 2021, rana, Update, PSL, HEPA speed raised Untitled.png
                   Reply  Tue May 18 20:26:11 2021, rana, Update, PSL, HEPA speed raised 
    Reply  Thu Apr 22 15:15:26 2021, gautam, Update, PSL, PMC transmission 
       Reply  Thu Apr 22 15:34:54 2021, Anchal, Update, PSL, PMC transmission 
Message ID: 16141     Entry time: Fri May 14 17:45:05 2021     In reply to: 16082     Reply to this: 16144
Author: rana 
Type: Update 
Category: PSL 
Subject: HEPA speed raised 

The PSL was too hot, so I turned on the south HEPA on the PSL. The north one was on and the south one was off (or so slow as to be inaudible and no vibration, unlike the north one). Lets watch the trend over the weekend and see if the temperature comes down and if the PMC / WFS variations get less. Fri May 14 17:46:26 2021

ELOG V3.1.3-